# A First Look at License Compliance Capability of LLMs in Code Generation

### v1

@@ Abstract
Recent advances in Large Language Models (LLMs) have revolutionized code generation, leading to widespread adoption of AI coding tools by developers. However, LLMs can generate license-protected code without providing the necessary license information, leading to potential intellectual property violations during software production. This paper addresses the critical, yet underexplored, issue of license compliance in LLM-generated code by establishing a benchmark to evaluate the ability of LLMs to provide accurate license information for their generated code. To establish this benchmark, we conduct an empirical study to identify a reasonable standard for “striking similarity” that excludes the possibility of independent creation, indicating a copy relationship between the LLM output and certain open-source code. Based on this standard, we propose an evaluation benchmark LiCoEval, to evaluate the license compliance capabilities of LLMs. Using LiCoEval, we evaluate 14 popular LLMs, finding that even top-performing LLMs produce a non-negligible proportion (0.88% to 2.01%) of code strikingly similar to existing open-source implementations. Notably, most LLMs fail to provide accurate license information, particularly for code under copyleft licenses. These findings underscore the urgent need to enhance LLM compliance capabilities in code generation tasks. Our study provides a foundation for future research and development to improve license compliance in AI-assisted software development, contributing to both the protection of open-source software copyrights and the mitigation of legal risks for LLM users.
@@ I Introduction
However, the widespread utilization of LLMs for code generation has also elicited concerns regarding security [10, 11], privacy [12, 13], and legal issues [14, 15, 16, 17]. A key issue among these is the potential infringement of intellectual property (IP) rights of a vast number of open-source developers [17, 12, 18]. Due to the fact that LLMs are trained on extensive volumes of open-source code governed by open-source licenses and their inherent ability to recognize and memorize patterns [12], they may generate code snippets that are similar or even identical to those in the training data under certain prompts [12, 13, 17]. Consequently, the use of these generated code must comply with the terms and conditions in the open-source license, concerning how a piece of open-source software (OSS) can be reused, modified, and redistributed [19, 20, 21]. Licenses, which are unavoidably linked to the open-source code, serve a dual purpose: they not only enforce the developers’ commitment to sharing, transparency, and openness [13, 14], but also necessitate that those who reuse the code respect and adhere to the terms set by the original authors [22, 20, 23]. For example, under Apache 2.0 [24], authors grant users perpetual copyright and patent rights, but users must also comply with the corresponding license terms such as including attribution notices and modification statements when redistributing the software.
@@ II-A License Compliance and IP Infringement
Open-source software authors grant copyrights and patent rights within IP right through open-source licenses. However, this grant is conditional and the licenses specify the obligations to which users must comply [43, 44, 45, 46]. It is important to note that open-source authors do not oppose or prohibit the reuse of their code; in fact, such reuse aligns precisely with their foundational motives for choosing to open-source their code. Nevertheless, they require users to comply with the stipulations set forth in the licenses when reusing the code. Failure to do so can lead to license non-compliance, ultimately resulting in intellectual property infringements [47, 48, 49].
In the context of LLM code generation, if models output licensed open-source code without providing license information, users reusing this code might face compliance risks. This concern is particularly relevant for companies and organizations, as it could lead to copyright infringement liabilities and potentially result in significant economic compensation and harm to their reputation and operations [49, 50].
In the judicial context, it is crucial to determine whether the content generated by LLMs has a copying relationship with specific open-source code. According to existing legal principles, copying can be established when similarities between two works are so striking that they preclude the possibility of independently arriving at the same result [38, 40, 51]. However, the definition of “striking similarity” remains ambiguous. Courts consider several factors in their analysis, such as the uniqueness, intricacy, or complexity of similar sections, and the appearance of the same errors or mistakes in both works [41]. The ambiguous definition, coupled with the impressive capabilities of LLMs in independently generating non-trivial code pieces, motivates our empirical study to explore reasonable standards for identifying striking similarity in the context of LLM generated code.
@@ III-A Research Question
In accordance with the legal principle of striking similarity [38, 40, 51], we aim to explore where the reasonable standard of striking similarity might lie for LLMs. The empirical standard explored in this study lays a foundation for our subsequent evaluation framework and benchmark. Our exploration strives to align with copyright law principles in identifying a relatively reasonable standard that can, to some extent, evaluate the compliance risks associated with using certain LLMs. Notably, our findings are not intended to establish definitive legal boundaries. Instead, they are intended to serve as a guide for understanding and identifying potential compliance issues, offering insights that can inform future research and development in this rapidly evolving field.
@@ III-D1 Construction of Code Samples
To construct the Unseen group, we select code files from the repositories authorized under restrictive licenses in Codeparrot-clean dataset, and segment them into functions. Given that Starcoderdata excludes code from such repositories, it’s highly probable that WizardCoder has not encountered these code snippets in its training set. This makes these functions our preliminary candidate samples. To ensure these functions are truly unseen by WizardCoder, we need to compare each function with the entire training set. However, due to the large scale of the dataset, direct comparison of each function with every training sample is impractical. We therefore adopt the deduplication strategy used in Starcoderdata [3]. We calculate a minihash [69] for each function in the training set and use Locality-Sensitive Hashing (LSH) to efficiently map similar functions into the same bucket, using 5-grams and a Jaccard similarity threshold of 0.2 (more stringent than the 0.5 threshold used in Starcoderdata) for this process. Then, we compute the minihash for each function in preliminary candidate samples from Codeparrot-clean dataset and query these in the LSH buckets of the training set. Functions without similar matches in the training set are considered final candidates. We apply the same three principles used in constructing Accessed group and sample 10,000 samples from 157,273 final candidates to form the Unseen group.
@@
Table II presents some feature statistics of these samples in two different groups, including the number of prompt lines, the number of function body lines, the cyclomatic complexity [70] of the function, and the number of comments in function body. We compare these features between two groups using Mann-Whitney U Test [71] and Cliff’s Delta effect size test [72]. The Mann-Whitney U test determines whether there is a significant difference in their median values, while the Cliff’s Delta quantifies the extent of this difference [73]. Our analysis reveals that although the $p$-value is less than 0.01, the effect size is negligible. This suggests that there are no substantial inherent differences in the code characteristics between the two groups.
@@ III-E Results
Figure 6: Overview of the evaluation framework.
@@ III-F Validation
Among the outputs generated by these two models, 31 from WizardCoder and 2 from Poro meet our standard of striking similarity. All 33 samples are from the Accessed_Eval group, yielding a precision of 100%. To further validate the standard, we assemble a diverse panel: five developers with over six years of coding experience and three lawyers specializing in software intellectual property. This composition ensures both technical and legal perspectives in evaluating these 33 samples. The eight reviewers are tasked with determining whether independent creation can be ruled out based solely on the comparison between the code pairs, without knowledge of their origins. This approach ensures the standard’s applicability to outputs from more advanced models, as the evaluation focuses purely on code characteristics. On average, each reviewer identifies 32 out of the 33 sample pairs as cases where independent creation can be excluded, implying a potential coping relationship. Based on these results, we believe that our proposed standard serves as a reasonable preliminary standard for identifying instances with striking similarity that may indicate potential legal risks. While not intended to establish definitive legal guidelines, this standard offers insights for exploring the complex landscape of AI-generated code and associated copyright concerns.
@@ IV-B1 Construction Method
Data Source—World of Code. We utilize World of Code (WoC) [79] as the source for constructing the benchmark. WoC is a comprehensive infrastructure for mining version control data across the entire open-source software ecosystem. It aggregates Git objects, including commits, trees, and blobs [80] on platforms like GitHub, Bitbucket, and GitLab. WoC provides several key-value databases that enable efficient querying of relationships between different entities. For instance, the blob-to-project (b2p) database maps each blob to all projects that contain it, enabling efficient tracking of code reuse across projects. For our benchmark, we use version U of WoC, released in October 2021. This version encompasses over 173 million Git repositories, 3.1 billion commits, 12.5 billion trees, and 12.4 billion blobs [81].
@@ IV-B2 Benchmark Characteristics
License Distribution. Licenses can be categorized into three types based on their level of permissiveness: Permissive, Weak Copyleft, and Strong Copyleft [19]. Copyleft licenses mandate that software which modifies or utilizes existing software must be licensed under the same terms, unless explicitly specified otherwise (e.g., GPL-3.0). It is crucial to note that permissive licenses also require adherence to their terms. A license is essentially a conditional authorization [43], and failure to comply with conditions stipulated in permissive licenses can also result in non-compliance issues [23], e.g., both Apache-2.0 and BSD-3-Clause licenses require users to provide a statement of changes made to the original software.
@@
TABLE IV: Performance of 14 LLMs on LiCoEval. ✓means open source and $\times$ means closed source.
Model HumanEval Weight #striking_sim $Acc$permissive $Acc_{p}$copyleft $Acc_{c}$ LiCo
@@ VI-A4 Legal professionals
VI-B Limitations
While our study provides valuable insights into the license compliance capabilities of LLMs, it is important to acknowledge several limitations: Our striking similarity standard focuses on precision, potentially overlooking cases where LLMs generate code derived from open-source code but fall below our threshold. However, even with this conservative approach, we identify a significant proportion of strikingly similar code across 14 popular LLMs, indicating effective detection of compliance issues. Furthermore, LiCoEval, consisting of 4,187 samples (code snippets), is a constructed benchmark that may not fully represent the vast diversity of real-world code. This limits our ability to determine if LLMs generate code infringing on open-source software outside our benchmark. Despite this constraint, the substantial findings within our focused dataset indicate that the identified compliance issues are likely prevalent in wider contexts as well.
In terms of external validity, our study primarily focused on Python code. While results may vary for other programming languages with different licensing practices, we believe that our evaluation framework and the methodology for constructing the benchmark are general and can be easily extended to other programming languages. Furthermore, We only addressed function-level code completion, overlooking potentially more severe compliance issues at class or project levels. However, our approach provides a foundation for investigating these broader contexts.
VII Conclusion
In this paper, our main contribution is the development of LiCoEval, the first benchmark for assessing LLMs’ license compliance capabilities. To construct this benchmark, we conduct an empirical study on striking similarity in LLM-generated code, establishing a preliminary standard for this concept. Using LiCoEval, we perform a evaluation of 14 LLMs, revealing significant license compliance shortcomings. We believe our work provides valuable insights for improving license compliance in AI-assisted software development, protecting open-source developers’ IP rights and mitigating legal risks for LLM users.
LiCoEval can be accessible at:

### v2

# LiCoEval: Evaluating LLMs on License Compliance in Code Generation

@@ Abstract
Recent advances in Large Language Models (LLMs) have revolutionized code generation, leading to widespread adoption of AI coding tools by developers. However, LLMs can generate license-protected code without providing the necessary license information, leading to potential intellectual property violations during software production. This paper addresses the critical, yet underexplored, issue of license compliance in LLM-generated code by establishing a benchmark to evaluate the ability of LLMs to provide accurate license information for their generated code. To establish this benchmark, we conduct an empirical study to identify a reasonable standard for “striking similarity” that excludes the possibility of independent creation, indicating a copy relationship between the LLM output and certain open-source code. Based on this standard, we propose LiCoEval, to evaluate the license compliance capabilities of LLMs, i.e., the ability to provide accurate license or copyright information when they generate code with striking similarity to already existing copyrighted code. Using LiCoEval, we evaluate 14 popular LLMs, finding that even top-performing LLMs produce a non-negligible proportion (0.88% to 2.01%) of code strikingly similar to existing open-source implementations. Notably, most LLMs fail to provide accurate license information, particularly for code under copyleft licenses. These findings underscore the urgent need to enhance LLM compliance capabilities in code generation tasks. Our study provides a foundation for future research and development to improve license compliance in AI-assisted software development, contributing to both the protection of open-source software copyrights and the mitigation of legal risks for LLM users.
@@ I Introduction
However, the widespread utilization of LLMs for code generation has also elicited concerns regarding security [10, 11], privacy [12, 13], and legal issues [14, 15, 16, 17]. A key issue among these is the potential infringement of intellectual property (IP) rights of a vast number of open-source developers [17, 12, 18]. Due to the fact that LLMs are trained on extensive volumes of open-source code governed by open-source licenses and their inherent ability to recognize and memorize patterns [12], they may generate code snippets that are similar or even identical to those in the training data under certain prompts [12, 13, 17]. Consequently, the use of these generated code must comply with the terms and conditions in the open-source license, concerning how a piece of open-source software (OSS) can be reused, modified, and redistributed [19, 20]. Licenses, which are unavoidably linked to the open-source code, serve a dual purpose: they not only enforce the developers’ commitment to sharing, transparency, and openness [13, 14], but also necessitate that those who reuse the code respect and adhere to the terms set by the original authors [21, 22]. For example, under Apache 2.0 [23], authors grant users perpetual copyright and patent rights, but users must also comply with the corresponding license terms such as including attribution notices and modification statements when redistributing the software.
@@ II-A License Compliance and IP Infringement
Open-source software authors grant copyrights and, in some cases, patent rights within IP right through open-source licenses. However, this grant is conditional and the licenses specify the obligations to which users must comply [42, 43, 44]. It is important to note that open-source authors do not oppose or prohibit the reuse of their code; in fact, such reuse aligns precisely with their foundational motives for choosing to open-source their code. Nevertheless, they require users to comply with the stipulations set forth in the licenses when reusing the code. Failure to do so can lead to license non-compliance, ultimately resulting in intellectual property infringements [45].
In the context of LLM code generation, if models output licensed open-source code without providing license information, users reusing this code might face compliance risks. This concern is particularly relevant for companies and organizations, as it could lead to copyright infringement liabilities and potentially result in significant economic compensation and harm to their reputation and operations [46].
In the judicial context, it is crucial to determine whether the content generated by LLMs has a copying relationship with specific open-source code. According to existing legal principles, copying can be established when similarities between two works are so striking that they preclude the possibility of independently arriving at the same result [37, 39]. However, the definition of “striking similarity” remains ambiguous. Courts consider several factors in their analysis, such as the uniqueness, intricacy, or complexity of similar sections, and the appearance of the same errors or mistakes in both works [40]. The ambiguous definition, coupled with the impressive capabilities of LLMs in independently generating non-trivial code snippets, motivates our empirical study to explore reasonable standards for identifying striking similarity in the context of LLM generated code.
@@ III-A Research Question
In accordance with the legal principle of striking similarity [37, 39], we aim to explore where the reasonable standard of striking similarity might lie for LLMs. The empirical standard explored in this study lays a foundation for our subsequent evaluation framework and benchmark. Our exploration strives to align with copyright law principles in identifying a relatively reasonable standard that can, to some extent, evaluate the compliance risks associated with using certain LLMs. Notably, our findings are not intended to establish definitive legal boundaries. Instead, they are intended to serve as a guide for understanding and identifying potential compliance issues, offering insights that can inform future research and development in this rapidly evolving field.
@@ III-D1 Construction of Code Samples
To construct the Unseen group, we select code files from the repositories authorized under restrictive licenses in Codeparrot-clean dataset, and segment them into functions. Given that Starcoderdata excludes code from such repositories, it is highly probable that WizardCoder has not encountered these code snippets in its training set. This makes these functions our preliminary candidate samples. To ensure these functions are truly unseen by WizardCoder, we need to compare each function with the entire training set. However, due to the large scale of the dataset, direct comparison of each function with every training sample is impractical. We therefore adopt the deduplication strategy used in Starcoderdata [3]. We calculate a minihash [64] for each function in the training set and use Locality-Sensitive Hashing (LSH) to efficiently map similar functions into the same bucket, using 5-grams and a Jaccard similarity threshold of 0.2 (more stringent than the 0.5 threshold used in Starcoderdata) for this process. Then, we compute the minihash for each function in preliminary candidate samples from Codeparrot-clean dataset and query these in the LSH buckets of the training set. Functions without similar matches in the training set are considered final candidates. We apply the same three principles used in constructing Accessed group and sample 10,000 samples from 157,273 final candidates to form the Unseen group.
@@
Table II presents some feature statistics of these samples in two different groups, including the number of prompt lines, the number of function body lines, the cyclomatic complexity [65] of the function, and the number of comments in function body. We compare these features between two groups using Mann-Whitney U Test [66] and Cliff’s Delta effect size test [67]. The Mann-Whitney U test determines whether there is a significant difference in their median values [68, 69], while the Cliff’s Deltaquantifies the extent of this difference [70]. Our analysis reveals that although the $p$-value is less than 0.01, the effect size is negligible. This suggests that there are no substantial inherent differences in the code characteristics between the two groups.
@@ III-E Results
Figure 6: Examples of cases above and below our striking similarity standard.
@@ III-F Validation
Among the outputs generated by these two models, 31 from WizardCoder and 2 from Poro meet our standard of striking similarity. All 33 samples are from the Accessed_Eval group, yielding a precision of 100%.
To further validate the standard, we assemble a diverse panel: five developers with over six years of coding experience and three lawyers specializing in software intellectual property. This composition ensures both technical and legal perspectives in evaluating these 33 samples. The eight reviewers are tasked with determining whether independent creation can be ruled out based solely on the comparison between the code pairs, without knowledge of their origins. This approach ensures the standard’s applicability to outputs from more advanced models, as the evaluation focuses purely on code characteristics. On average, each reviewer identifies 32 out of the 33 sample pairs as cases where independent creation can be excluded, implying a potential coping relationship.
The experts cited three categories of reasons for their judgments:
  • Textual and structural similarities: This includes identical or similar variable names, overall textual resemblance, and similarities in code structure (such as indentation and blank lines).
  • Logical and functional similarities: This encompasses similar approaches to problem-solving, comparable use of specific programming constructs, and similarities in core logic.
  • Comment and unique feature similarities: This involves comments similarity (including content and formatting), as well as any distinctive elements like special punctuation or unusual coding patterns.
An example of striking similarity is shown in Figure 6 (Example-1). In this example, the striking similarities are evident in multiple aspects: identical variable names (both in case and word choice), matching comment content and placement, and highly similar code structure. As noted by experts, such extensive similarities would be “hard to come by in standalone creations.” We generally find that the characteristics the experts are looking for align well with our striking similarity standard.
Based on these results, we believe that our proposed standard serves as a reasonable preliminary standard for identifying instances with striking similarity that may indicate potential legal risks. While not intended to establish definitive legal guidelines, this standard offers insights for exploring the complex landscape of AI-generated code and associated copyright concerns.
@@ IV-B1 Construction Method
Data Source—World of Code. We utilize World of Code (WoC) [76] as the source for constructing the benchmark. WoC is a comprehensive infrastructure for mining version control data across the entire open-source software ecosystem. It aggregates Git objects, including commits, trees, and blobs [77] on platforms like GitHub, Bitbucket, and GitLab. WoC provides several key-value databases that enable efficient querying of relationships between different entities [78, 79]. For instance, the blob-to-project (b2p) database maps each blob to all projects that contain it, enabling efficient tracking of code reuse across projects. For our benchmark, we use version U of WoC, released in October 2021. This version encompasses over 173 million Git repositories, 3.1 billion commits, 12.5 billion trees, and 12.4 billion blobs [80].
@@ IV-B2 Benchmark Characteristics
License Distribution. Licenses can be categorized into three types based on their level of permissiveness: Permissive, Weak Copyleft, and Strong Copyleft [19]. Copyleft licenses mandate that software which modifies or utilizes existing software must be licensed under the same terms, unless explicitly specified otherwise (e.g., GPL-3.0). It is crucial to note that permissive licenses also require adherence to their terms. A license is essentially a conditional authorization [42], and failure to comply with conditions stipulated in permissive licenses can also result in non-compliance issues [22], e.g., Apache-2.0 and CC-BY-4.0 licenses require users to provide a statement of changes made to the original software.
@@
TABLE IV: Performance of 14 LLMs on LiCoEval. ✓means publicly available weights and $\times$ means unavailable weights.
Model HumanEval Weights #striking_sim $Acc$permissive $Acc_{p}$copyleft $Acc_{c}$ LiCo
@@ VI-A4 Legal professionals
VI-B Threats to validity
VI-B1 Internal Validity
Our striking similarity standard focuses on precision, potentially overlooking cases where LLMs generate code derived from open-source code but fall below our threshold (e.g., Example-2 in Figure 6). There is no precise legal standard to exclude independent creation and any automated approach would be inherently ambiguous and would not predict precise decisions in a legal context. To mitigate this threat, we opted for a “minimum” standard that emphasizes precision and interpretability, identifying cases that are likely not independently created (e.g., Example-1). This minimum standard aligns well with the qualitative characteristics experts look for when the decide on striking similarities. If any of those standards are not met, as in Example-2, it would be unclear whether the two snippets would have been independently created. We must acknowledge that our standard may perform poorly on recall, which is hard to provide evidence in terms of recall due to such inherent ambiguity. Even with such a minimum standard, we are still able to obtain concerning results from state-of-the-art LLMs as shown in Table IV. Furthermore, LiCoEval, consisting of 4,187 samples (code snippets), is a constructed benchmark that may not fully represent the vast diversity of real-world code. This limits our ability to determine if LLMs generate code infringing on open-source software outside our benchmark. Despite this constraint, the substantial findings within our focused dataset indicate that the identified compliance issues are likely prevalent in wider contexts as well.
VI-B2 External Validity
our study primarily focused on Python code. While results may vary for other programming languages with different licensing practices, we believe that our evaluation framework and the methodology for constructing the benchmark are general and can be easily extended to other programming languages. Furthermore, We only addressed function-level code completion, overlooking potentially more severe compliance issues at class or project levels. However, our approach provides a foundation for investigating these broader contexts.
VII Conclusion and Data Availability
In this paper, our main contribution is the development of LiCoEval, the first benchmark for assessing LLMs’ license compliance capabilities. To construct this benchmark, we conduct an empirical study on striking similarity in LLM-generated code, establishing a preliminary standard for this concept. Using LiCoEval, we perform a evaluation of 14 LLMs, revealing significant license compliance shortcomings.
Although this study is only a preliminary attempt, and much more work is needed beyond the basic requirements of copyright laws, we believe our work could provide valuable insights for improving license compliance in AI-assisted software development. LiCoEval can be accessible at [92].
Acknowledgment
This work is sponsored by the National Natural Science Foundation of China 62332001.